G3tH3Pxmpthat5Amx5t:

What is the solution to the equation StartFraction 1 Over h minus 5 EndFraction + StartFraction 2 Over h + 5 EndFraction = StartFraction 16 Over h squared minus 25 EndFraction?

IamAku:

please put this under the correct subject

jhonyy9:

@IamAku this is math question on math subject - why not is correct posted ?

jhonyy9:

\[\frac{ 1 }{ h - 5 } + \frac{ 2 }{ h + 5 } = \frac{ 16 }{ h^2 - 25 }\] so than you think on these in this way what is the first step - how you add two fractions ? did you heard about common denominator ?